In 5 year's time, the sum of the ages ofna pair of twins and their elder sister is 45. The present age of the elder sister is 6 years less than the sum or the present age of the twins.Find the present ages of each twins and their elder sister
Answer by ankor@dixie-net.com(22740)

In 5 year's time, the sum of the ages of a pair of twins and their elder sister is 45.
The present age of the elder sister is 6 years less than the sum or the present age of the twins.
Find the present ages of each twins and their elder sister
:
let s = elder sister's age
let t = age of the twins
:
Write an equation for each statement, simplify as much as possible:
:
"In 5 year's time, the sum of the ages of a pair of twins and their elder sister is 45."
s+5 + t+5 + t+5 = 45
s + 2t + 15 = 45
s + 2t = 45-15
s + 2t = 30
:
"present age of the elder sister is 6 years less than the sum of the present age of the twins"
s = 2t - 6
:
Replace s in the 1st equation with (2t-6)
(2t-6) + 2t = 30
2t + 2t = 30 + 6
4t = 36
t = 9 yrs old are the twins
then
s = 2(9) - 6
s = 18 - 6
s = 12 yrs is the age of the sister
:
:
See if that checks out in the 1st statement
"In 5 year's time, the sum of the ages of a pair of twins and their elder sister is 45."
12 + 5 + 9 + 5 + 9 + 5 = 45
:
:
I'm not comfortable with the phrase "a pair of twins", we know twins are a pair, would a pair of twins be 4 people? Probably just poor writing. CK